OpenStudy (anonymous):

14x^6- 8x^4 Divided by 2x^3 (the first part of the numbers are at the top and the second at the bottom of a fraction)

OpenStudy (mertsj):

Think of it like this: \[\frac{14x^6}{2x^3}-\frac{8x^4}{2x^3}\]

OpenStudy (anonymous):

but the exponenets are different so I still don't know how to solve :/

OpenStudy (mertsj):

Use this property: \[\frac{a^m}{a^n}=a ^{m-n}\]

OpenStudy (anonymous):

making the answer 7x^3-4x^-1 ?

OpenStudy (mertsj):

\[7x^3 -4x^1\]

OpenStudy (anonymous):

Thanks!!